a collection of 35 coins contains nickels and quarters. find the number of nickels given that the total face value of the coins is $2.55.
Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

a collection of 35 coins contains nickels and quarters. find the number of nickels given that the total face value of the coins is $2.55.
---
Quantity: n + q = 35 coins
Value:: 5n+ 25q = 255 cents
--------------------------------
Divide thru the Value Eq. by 5:
n + 5q = 51
n + q = 35
-------------
Subtract and solve for "q":
4q = 16
q = 4 (# of quarters)
--------
Since n + q = 35, n = 31 (# of nickels)
------------------------------------
Cheers,
Stan H.

Answer by ewatrrr(24785)


Hi,
Let x and (35-x)represent the quarters and nickels respectively
Question states:
25x + 5(35-x) = 255 cents
20x = 255 - 5*35
20x = 80
x = 4, the number of quarters, number of nickels = 31 (35-4)
CHECKING our Answer
25*4 + 5*31 = 100 + 155 = 255 cents = $2.55